# Constrained Stein Variational Trajectory Optimization

### v2

@@ Abstract
We present Constrained Stein Variational Trajectory Optimization (CSVTO), an algorithm for performing trajectory optimization with constraints on a set of trajectories in parallel. We frame constrained trajectory optimization as a novel form of constrained functional minimization over trajectory distributions, which avoids treating the constraints as a penalty in the objective and allows us to generate diverse sets of constraint-satisfying trajectories. Our method uses Stein Variational Gradient Descent (SVGD) to find a set of particles that approximates a distribution over low-cost trajectories while obeying constraints. CSVTO is applicable to problems with differentiable equality and inequality constraints and includes a novel particle re-sampling step to escape local minima. By explicitly generating diverse sets of trajectories, CSVTO is better able to avoid poor local minima and is more robust to initialization. We demonstrate that CSVTO outperforms baselines in challenging highly-constrained tasks, such as a 7DoF wrench manipulation task, where CSVTO outperforms all baselines both in success and constraint satisfaction. Our results demonstrate that generating diverse constraint-satisfying trajectories improves robustness to disturbances and initialization over baselines.
@@ I Introduction
One of the key advantages of trajectory optimization techniques over global search methods, such as sampling-based motion planning, is computational speed. Faster computational speed enables online re-planning to adapt to disturbances. For example, consider again the robot sanding the table, but now in the proximity of a human. The human may move in an unexpected way which necessitates an update to the planned trajectory. However, the performance of many trajectory optimization methods is highly dependent on the initialization. Poor initialization may lead to the solver converging to a poor local minimum. For example, for a robot minimizing a distance to goal cost subject to collision constraints, this may mean a trajectory that avoids obstacles but makes little or no progress toward the goal. In the worst case, the solver may not find a feasible solution, in which case the robot may collide with an obstacle. A dependence on initialization is particularly problematic when re-solving the optimization problem online under limited computation time when disturbances can lead to the previous solution becoming a poor initialization for the current optimization problem. In the sanding example mentioned previously, the human may move to block the robot’s path, and performing a local optimization starting from the previous trajectory may not return a feasible solution.
@@
We propose Constrained Stein Variational Trajectory Optimization (CSVTO), an algorithm that performs constrained trajectory optimization on a set of trajectories in parallel. Our method builds on Orthogonal-Space Stein Variational Gradient Descent (O-SVGD), a recent non-parametric variational inference method for domains with a single equality constraint [13]. We present a constrained SVGD algorithm for trajectory optimization with differentiable equality and inequality constraints, generating a diverse set of approximately constraint-satisfying trajectories. The trajectories are approximately constraint-satisfying because we do not run the algorithm until convergence to avoid excessive computation times. We additionally incorporate a novel re-sampling step that re-samples and perturbs particles in the tangent space of the constraints to escape local minima. Our contributions are as follows:
@@
The article is organized as follows. In Section II we discuss related work. In Section III we will discuss the trajectory optimization problem, followed by an overview of the variational inference approach to trajectory optimization in Section IV. In Section V we introduce our novel formulation of trajectory optimization as a constrained functional minimization over trajectory distributions. We will then give some additional background information on SVGD in Section VI which is necessary to develop our algorithm. In Section VII we introduce CSVTO. In Section VIII we evaluate our method on a 12DoF quadrotor task and two highly constrained tasks with a 7DoF manipulator. We additionally deployed CSVTO to turn a wrench in the real world (Figure 1).
@@ II-C Planning & Control as Inference
Another class of methods has used Stein Variational Gradient Descent (SVGD) [39] for Model Predictive Control (MPC) [10, 40] and trajectory optimization [6]. By using particle approximations these methods can generate multi-modal trajectory distributions. SVGD has also been used to improve Probabilistic Roadmaps (PRMs) [41]. Our method is also based on SVGD.
@@ III Trajectory Optimization
Here we have separated general inequality constraints $g$ from simple bounds constraints, as well as the dynamics constraints from other equality constraints $h$. We additionally assume that $C$ is non-negative and once differentiable and that $f,g,h$ are all twice differentiable [^1] We can also accommodate constraints that are only once-differentiable via an approximation (see Section VII-A1). Problem (1) will be non-convex in general, therefore it is likely it will have multiple local minima. The quality of solutions for most methods for solving this optimal control problem depends heavily on the initialization; often a poor initialization can lead to infeasibility.
@@ IV Variational Inference for Trajectory Optimization
We aim to find the posterior distribution $p(\tau|o=1)\propto p(o=1|\tau)p(\tau)$, where $p(\tau)=p(\mathbf{X},\mathbf{U})$ is a prior on trajectories. For deterministic dynamics, this prior is determined by placing a prior on controls $\mathbf{U}$. This prior is a design choice and can be used to regularize the controls. For instance, a squared control cost can be equivalently expressed as a Gaussian prior. Alternatively, this prior could be learned from a dataset of low-cost trajectories [53]. The trajectory prior is
@@
where $\mathbf{\hat{x}}_{t}=f(\mathbf{x}_{t-1},\mathbf{u}_{t-1})$, and $\delta$ is the Diract delta function. This inference problem can be performed exactly for the case of linear dynamics and quadratic costs [54, 29]. However, in general, this problem is intractable and approximate inference techniques must be used. We use variational inference to approximate $p(\tau|o=1)$ with distribution $q(\tau)$ which minimizes the Kullback–Leibler divergence $\mathcal{KL}(q(\tau)||p(\tau|o=1))$ [7]. The quantity to be minimized is
@@ V Problem Statement
Our goal is to solve the above optimization problem. However, for any practical algorithm, we cannot guarantee exact constraint satisfaction, both due to the potential non-convexity of the constraint functions and due to numerical issues. This is particularly true when planning online with limited computational time. Therefore we will evaluate our method according to both the cost of the resulting trajectories and the amount of constraint violation when optimizing within a fixed number of iterations.
@@ VI Stein Variational Gradient Descent
We develop an algorithm to solve the constrained variational inference objective in (9) based on Stein Variational Gradient Descent (SVGD) [39]. In this section we will give an overview of SVGD which forms the foundation of our method. SVGD is a variational inference technique that uses a non-parametric representation of the variational posterior. In our algorithm, we use SVGD to approximate the distribution $p(\tau|o=1)$ with particles, where each particle is a trajectory. Consider the variational inference problem
@@
where $\epsilon>0$ is a step-size parameter. The update $\phi^{*}$ is computed using a differentiable positive definite kernel function $\mathcal{K}$ via
@@
where $\nabla_{\mathbf{x}}\cdot\phi(\mathbf{x})=\sum^{d}_{k=1}\partial_{x_{k}}\phi_{k}(\mathbf{x})$. It was established in [55] that $\mathbb{S}(q,p)=0\iff p=q$ for a strictly positive-definite kernel $\mathcal{K}$. To minimize the KL divergence SVGD considers the incremental transform $\mathbf{x}_{\epsilon}=\mathbf{x}+\epsilon\phi(\mathbf{x})$, where $\mathbf{x}\sim q(\mathbf{x})$ and $\epsilon$ is a scalar step-size parameter. The resulting distribution after applying the transform is $q_{[\epsilon\phi]}$. SVGD uses the following result:
@@ VII Methods
Our proposed trajectory optimization algorithm uses SVGD to perform constrained optimization on a set of trajectories in parallel. The result is a diverse set of low-cost constraint-satisfying trajectories. The full algorithm is shown in Algorithm 1. First, we will introduce the main component of our proposed algorithm, which decomposes the Stein update into a step tangential to the constraint boundary, and a step toward constraint satisfaction. We will then provide an analysis of the algorithm which relates it to problem (9). Finally, we will discuss strategies for improving performance which include separating the bounds constraints, an annealing strategy for increasing particle diversity, and re-sampling particles during the optimization. Figure 2 demonstrates CSVTO being applied to a 2D toy problem.
@@ VII-A Constrained Stein Trajectory Optimization
Solving the constrained variational inference problem in (9) is very difficult, since it requires finding a distribution that may exhibit multi-modality and has constrained support. To deal with this, we adopt a non-parametric approach to representing distribution $q(\tau)$. We use SVGD where each particle is a trajectory, and iteratively update the particle set while enforcing the constraints on each particle. To do this we extend O-SVGD to multiple equality and inequality constraints and use it to generate constraint-satisfying trajectories.
@@ VII-A1 Equality constraints
We propose a modified Stein update rule in which we decompose the update into two components:
@@
where ϕsubscriptitalic-$\phi_{\perp}$ is an update that is tangential to the constraint boundary, $\phi_{C}$ acts in the direction which decreases constraint violation, and $\alpha_{J}$ and $\alpha_{C}$ are scalar step size parameters. We replace the O-SVGD $\phi_{C}$ from equation (24) with a Gauss-Newton step to minimize $h(\tau)^{T}h(\tau)$
@@
Performing this inverse is only possible if $\nabla h(\tau)$ full rank. If it is not full rank, we instead use the pseudo-inverse of $\nabla h(\tau)\nabla h(\tau)^{T}$. Once we have $P(\tau)$, we use it to define the tangent space kernel, as in [13]:
@@ VII-A2 Extension to Inequality Constraints
We extend the above method to inequality constraints with the use of slack variables. We turn the inequality constraints into equality constraints with slack variable $z$:
$g(\tau)+\frac{1}{2}z^{2}=0.$ (33)
Converting inequality constraints to equality constraints via squared slack variables is often avoided as it can introduce spurious non-local-minima that satisfy the Karush–Kuhn–Tucker (KKT) conditions [57]. To mitigate this issue we make an assumption on the regularity of the problem, denoted as (R) in [46]. The details of the assumption are technical and we do not include it here. The assumption essentially states that $\nabla\hat{h}$ is full rank at initialization and remains so during the optimization. Under these assumptions, Schropp [46] proved that the hyperbolic equilibrium points of the augmented system are local minima of the equality and inequality-constrained optimization problem. Optimizing multiple trajectories in parallel provides additional robustness against this issue. Even should some particles become stuck at one of these undesirable fixed points, in Section VII-A8 we propose a method for re-sampling the set of particles which redistributes particles away from these fixed points. While we could avoid this issue by using non-negative slack variables with the transformation $g(\tau)+z$, where $z>0$, we found that this led to poorer constraint satisfaction in practice.
After introducing the slack variables, we compute the constrained Stein update with all constraints as equality constraints. The full set of equality constraints then becomes
$\hat{h}=\begin{bmatrix}h(\tau)\\ g(\tau)+\frac{1}{2}z^{2}\end{bmatrix}.$ (34)
We augment the state with $z$ as
$\hat{\tau}=\begin{bmatrix}\tau\\ z\end{bmatrix}.$ (35)
@@ VII-A3 Analysis
In this section, we provide an analysis of CSVTO. We demonstrate that stationary points of the gradient flow satisfy the first-order optimality conditions for the constrained variational optimization problem in (9), subject only to equality constraints.
@@
Since $\phi_{C}$ and ϕ*subscriptsuperscriptitalic-$\phi^{*}_{\perp}$ are orthogonal, then if equation (41) holds then $\phi_{C}=\phi^{*}_{\perp}=0$. Next, we note that $\phi^{*}_{\perp}(\tau)=P(\tau)\hat{\phi}$, where $\hat{\phi}\in\mathcal{H}^{d}$ and further $P(\tau)\hat{\phi}(\tau)=0\implies P(\tau)\phi^{*}(\tau)=0$. To see this, consider $P(\tau)\phi^{*}(\tau)\neq 0$. This would imply that $\nabla_{\epsilon}\mathcal{KL}(q_{[\epsilon P\phi^{*}]}||p(\tau|o=1))|_{\epsilon=0}\neq 0$, which implies that there is a descent direction. This would mean that ϕsubscriptitalic-$\exists\>\phi_{\perp}$ such that $-\nabla_{\epsilon}\mathcal{KL}(q_{[\epsilon\phi_{\perp}^{*}]}||p(\tau|o=1))|_{\epsilon=0}<-\nabla_{\epsilon}\mathcal{KL}(q_{[\epsilon\phi_{\perp}}||p(\tau|o=1))_{\epsilon=0}$, which is a contradiction. Expanding $P(\tau)\phi^{*}=0$ yields

### v3

@@ Abstract
We present Constrained Stein Variational Trajectory Optimization (CSVTO), an algorithm for performing trajectory optimization with constraints on a set of trajectories in parallel. We frame constrained trajectory optimization as a novel form of constrained functional minimization over trajectory distributions, which avoids treating the constraints as a penalty in the objective and allows us to generate diverse sets of constraint-satisfying trajectories. Our method uses Stein Variational Gradient Descent (SVGD) to find a set of particles that approximates a distribution over low-cost trajectories while obeying constraints. CSVTO is applicable to problems with differentiable equality and inequality constraints and includes a novel particle re-sampling step to escape local minima. By explicitly generating diverse sets of trajectories, CSVTO is better able to avoid poor local minima and is more robust to initialization. We demonstrate that CSVTO outperforms baselines in challenging highly-constrained tasks, such as a 7DoF wrench manipulation task, where CSVTO outperforms all baselines both in success and constraint satisfaction.
@@ I Introduction
Global sample-based motion planning methods such as Rapidly-Exploring-Random-Trees (RRT) [6]. Probabilistic Roadmaps (PRM) [7] effectively solve difficult planning problems, however, they do not find paths that minimize a given cost function. To minimize a given cost function, algorithms such as RRT* and PRM* [8] have been proposed to find asymptotically globally optimal paths. Alternatively, a common approach is to use the path returned from a sample-based motion planner to initialize a trajectory optimization problem [9]. Sample-based methods have additionally been applied to constrained planning problems [10, 11, 12, 13], and kinodynamic problems [14, 15]. While effective for solving problems exhibiting local minima, when applied to kinodynamic or constrained problems these global methods are typically computationally expensive.
@@
One of the key advantages of trajectory optimization techniques over global search methods, such as sampling-based motion planning, is computation speed. Faster computation speed enables online re-planning to adapt to disturbances. For example, consider again the robot sanding the table, but now in the proximity of a human. The human may move in an unexpected way which necessitates an update to the planned trajectory. However, even if the cost function is well-suited to the task, the performance of many trajectory optimization methods is still highly dependent on the initialization. Poor initialization may lead to the solver converging to a poor local minimum. For example, for a robot minimizing a distance to goal cost subject to collision constraints, this may mean a trajectory that avoids obstacles but makes little or no progress toward the goal. In the worst case, the solver may not find a feasible solution, in which case the robot may collide with an obstacle. A dependency on initialization is particularly problematic when re-solving the optimization problem online under limited computation time when disturbances can lead to the previous solution becoming a poor initialization for the current optimization problem. In the sanding example mentioned previously, the human may move to block the robot’s path, and performing a local optimization starting from the previous trajectory may not return a feasible solution.
@@
We propose Constrained Stein Variational Trajectory Optimization (CSVTO), an algorithm that performs constrained trajectory optimization on a set of trajectories in parallel. Our method builds on Orthogonal-Space Stein Variational Gradient Descent (O-SVGD), a recent non-parametric variational inference method for domains with a single equality constraint [23]. We present a constrained Stein Variational Gradient Descent (SVGD) algorithm for trajectory optimization with differentiable equality and inequality constraints, generating a diverse set of approximately constraint-satisfying trajectories. The trajectories are approximately constraint-satisfying because we do not run the algorithm until convergence to avoid excessive computation times. We additionally incorporate a novel re-sampling step that re-samples and perturbs particles in the tangent space of the constraints to escape local minima. Our contributions are as follows:
@@
The rest of the article is organized as follows. In Section II we discuss related work. In Section III we will discuss the trajectory optimization problem, followed by an overview of the variational inference approach to trajectory optimization in Section IV. In Section V we introduce our novel formulation of trajectory optimization as a constrained functional minimization over trajectory distributions. We will then give some additional background information on SVGD in Section VI which is necessary to develop our algorithm. In Section VII we introduce CSVTO. In Section VIII we evaluate our method on a 12DoF quadrotor task and two highly constrained tasks with a 7DoF manipulator. We additionally deployed CSVTO to turn a wrench in the real world (Figure 1).
@@ II-C Planning & Control as Inference
Another class of methods has used Stein Variational Gradient Descent (SVGD) [45] for Model Predictive Control (MPC) [20, 46] and trajectory optimization [16]. By using particle approximations these methods can generate multi-modal trajectory distributions. SVGD has also been used to improve Probabilistic Roadmaps (PRMs) [47], and for planning to goal sets [48]. Our method is also based on SVGD.
@@ III Trajectory Optimization
Here we have separated general inequality constraints $g$ from simple bounds constraints, as well as the dynamics constraints from other equality constraints $h$. We additionally assume that $C$ is non-negative and once differentiable and that $f,g,h$ are all twice differentiable [^1] We can also accommodate constraints that are only once-differentiable via an approximation (see Section VII-A1. Problem (1) will be non-convex in general, therefore it is likely it will have multiple local minima. The quality of solutions for most methods for solving this optimal control problem depends heavily on the initialization; often a poor initialization can lead to infeasibility.
@@ IV Variational Inference for Trajectory Optimization
We aim to find the posterior distribution over trajectories, conditioned on the value of auxiliary variable $o$. This is given by Bayes theorem as
$p(\tau|o=1)=\frac{p(o=1|\tau)p(\tau)}{p(o=1)},$ (3)
where $p(\tau)=p(\mathbf{X},\mathbf{U})$ is a prior on trajectories. For deterministic dynamics, this prior is determined by placing a prior on controls $\mathbf{U}$. This prior is a design choice and can be used to regularize the controls. For instance, a squared control cost can be equivalently expressed as a Gaussian prior. Alternatively, this prior could be learned from a dataset of low-cost trajectories [60]. The trajectory prior is
@@
where $\mathbf{\hat{x}}_{t}=f(\mathbf{x}_{t-1},\mathbf{u}_{t-1})$, and $\delta$ is the Dirac delta function. This inference problem can be performed exactly for the case of linear dynamics and quadratic costs [61, 35]. However, in general, this problem is intractable and approximate inference techniques must be used. We use variational inference to approximate $p(\tau|o=1)$ with distribution $q(\tau)$ which minimizes the Kullback–Leibler divergence $\mathcal{KL}(q(\tau)||p(\tau|o=1))$ [17]. The quantity to be minimized is
@@ V Problem Statement
Our goal is to solve the above optimization problem. However, for any practical algorithm, we cannot guarantee exact constraint satisfaction, both due to the potential non-convexity of the constraint functions and due to limited computation time. Computation time is especially limited in an online planning scenario. Therefore we will evaluate our method according to both the cost of the resulting trajectories and the amount of constraint violation when optimizing within a fixed number of iterations.
@@ VI Stein Variational Gradient Descent
We develop an algorithm to solve the constrained variational inference objective in (LABEL:eq:constrained_planning_inference) based on Stein Variational Gradient Descent (SVGD) [45]. In this section we will give an overview of SVGD which forms the foundation of our method. SVGD is a variational inference technique that uses a non-parametric representation of the variational posterior. In our algorithm, we use SVGD to approximate the distribution $p(\tau|o=1)$ with particles, where each particle is a trajectory. Consider the variational inference problem
@@
where $\epsilon>0$ is a step-size parameter, $k$ is the iteration number, and $i$ is the particle index. The update $\phi^{*}$ is computed using a differentiable positive definite kernel function $\mathcal{K}$ via
@@
where $\nabla_{\mathbf{x}}\cdot\phi(\mathbf{x})=\sum^{d}_{k=1}\partial_{x_{k}}\phi_{k}(\mathbf{x})$. It was established in [62] that $\mathbb{S}(q,p)=0\iff p=q$ for a strictly positive-definite kernel $\mathcal{K}$. To minimize the KL divergence, SVGD considers the incremental transform $\mathbf{x}_{\epsilon}=\mathbf{x}+\epsilon\phi(\mathbf{x})$, where $\mathbf{x}\sim q(\mathbf{x})$ and $\epsilon$ is a scalar step-size parameter. The resulting distribution after applying the transform is $q_{[\epsilon\phi]}$. SVGD uses the following result:
@@ VII Methods
Our proposed trajectory optimization algorithm uses SVGD to perform constrained optimization on a set of trajectories in parallel. The result is a diverse set of low-cost constraint-satisfying trajectories. The full algorithm is shown in Algorithm 1. First, we will introduce the main component of our proposed algorithm, which decomposes the Stein update into a step tangential to the constraint boundary, and a step toward constraint satisfaction. We will then provide an analysis of the algorithm which relates it to problem (LABEL:eq:constrained_planning_inference). Finally, we will discuss strategies for improving performance which include separating the bounds constraints, an annealing strategy for increasing particle diversity, and re-sampling particles during the optimization. Figure 2 demonstrates CSVTO being applied to a 2D toy problem.
@@ VII-A Constrained Stein Trajectory Optimization
Solving the constrained variational inference problem in (LABEL:eq:constrained_planning_inference) is very difficult, since it requires finding a distribution that may exhibit multi-modality and has constrained support. To address this, we use a non-parametric representation of the distribution $q(\tau)$. We use SVGD where each particle is a trajectory, and iteratively update the particle set while enforcing the constraints on each particle. To do this we extend O-SVGD to multiple equality and inequality constraints and use it to generate constraint-satisfying trajectories.
@@ VII-A1 Equality constraints
We propose a modified Stein update rule for the $i$-th particle, in which we decompose the update into two components:
@@
where ϕsubscriptitalic-$\phi_{\perp}$ is an update that is tangential to the constraint boundary, $\phi_{C}$ acts in the direction that decreases constraint violation, $\alpha_{J}$ and $\alpha_{C}$ are scalar step size parameters, and $k$ is the iteration. We replace the O-SVGD $\phi_{C}$ from equation (25) with a Gauss-Newton step to minimize $h(\tau)^{T}h(\tau)$
@@
Inverting $\nabla h(\tau)\nabla h(\tau)^{T}$ is only possible if $\nabla h(\tau)$ is full rank. While in Sections VII-A2 and VII-A3 we assume that $\nabla h(\tau)$ is full rank, for numerical stability we compute the pseudo-inverse of $\nabla h(\tau)\nabla h(\tau)^{T}$ via the singular-value decomposition, discarding singular vectors corresponding to singular values that are smaller than $1\times 10^{-6}$. Once we have $P(\tau)$, we use it to define the tangent space kernel, as in [23]:
@@ VII-A2 Extension to Inequality Constraints
We extend the above method to inequality constraints with the use of slack variables. We turn the inequality constraints into equality constraints with slack variable $\mathbf{z}$:
$g(\tau)+\frac{1}{2}\mathbf{z}^{2}=0.$ (34)
The full set of equality constraints then becomes
$\hat{h}=\begin{bmatrix}h(\tau)\\ g(\tau)+\frac{1}{2}\mathbf{z}^{2}\end{bmatrix}.$ (35)
Converting inequality constraints to equality constraints via squared slack variables is often avoided as it can introduce spurious non-local-minima that satisfy the Karush–Kuhn–Tucker (KKT) conditions [64]. To mitigate this issue we make an assumption on the regularity of the problem, denoted as (R) in [53]. The details of the assumption are technical and we do not include it here. The assumption essentially states that $\nabla\hat{h}$ is full rank at initialization and remains so during the optimization. Under these assumptions, Schropp [53] proved that the hyperbolic equilibrium points of the augmented system are local minima of the equality and inequality-constrained optimization problem. Optimizing multiple trajectories in parallel provides additional robustness against this issue. Even should some particles become stuck at one of these undesirable fixed points, in Section VII-A8 we propose a method for re-sampling the set of particles which redistributes particles away from these fixed points. While we could avoid this issue by using non-negative slack variables with the transformation $g(\tau)+\mathbf{z}$, where $\mathbf{z}>0$, we found that this led to poorer constraint satisfaction in practice.
After introducing the slack variables, we compute the constrained Stein update with all constraints as equality constraints. We augment the state with $\mathbf{z}$ as
$\hat{\tau}=\begin{bmatrix}\tau\\ \mathbf{z}\end{bmatrix}.$ (36)
@@ VII-A3 Analysis
In this section, we provide an analysis of CSVTO. We demonstrate that stationary points of the gradient flow satisfy the first-order optimality conditions for the constrained variational optimization problem in (LABEL:eq:constrained_planning_inference), subject only to equality constraints.
@@
Since $\phi_{C}$ and ϕsubscriptsuperscriptitalic-$\phi^{*}_{\perp}$ are orthogonal, then if equation (42) holds then $\phi_{C}=\phi^{*}_{\perp}=0$. Next, we note that $\phi^{*}_{\perp}(\tau)=P(\tau)\hat{\phi}$, where $\hat{\phi}\in\mathcal{H}^{d}$ and further $P(\tau)\hat{\phi}(\tau)=0\implies P(\tau)\phi^{*}(\tau)=0$. To see this, consider $P(\tau)\phi^{*}(\tau)\neq 0$. This would imply that $\nabla_{\epsilon}\mathcal{KL}(q_{[\epsilon P\phi^{*}]}||p(\tau|o=1))|_{\epsilon=0}\neq 0$, which implies that there is a descent direction. This would mean that ϕsubscriptitalic-$\exists\>\phi_{\perp}$ such that $-\nabla_{\epsilon}\mathcal{KL}(q_{[\epsilon\phi_{\perp}^{*}]}||p(\tau|o=1))|_{\epsilon=0}<-\nabla_{\epsilon}\mathcal{KL}(q_{[\epsilon\phi_{\perp}}||p(\tau|o=1))_{\epsilon=0}$, which is a contradiction. Expanding $P(\tau)\phi^{*}=0$ yields